# Annotator-Centric Active Learning for Subjective NLP Tasks

### v1

@@ Abstract
To accurately capture the variability in human judgments for subjective NLP tasks, incorporating a wide range of perspectives in the annotation process is crucial. Active Learning (AL) addresses the high costs of collecting human annotations by strategically annotating the most informative samples. We introduce Annotator-Centric Active Learning (ACAL), which incorporates an annotator selection strategy following data sampling. Our objective is two-fold: (1) to efficiently approximate the full diversity of human judgments, and (2) to assess model performance using annotator-centric metrics, which emphasize minority perspectives over a majority. We experiment with multiple annotator selection strategies across seven subjective NLP tasks, employing both traditional and novel, human-centered evaluation metrics. Our findings indicate that ACAL improves data efficiency and excels in annotator-centric performance evaluations. However, its success depends on the availability of a sufficiently large and diverse pool of annotators to sample from.
@@ 1 Introduction
A challenging aspect of natural language understanding (NLU) is the variability of human judgment and interpretation in subjective tasks (e.g., hate speech detection) (Plank, 2022). While humans can navigate subjectivity naturally, most machine learning methods are insensitive to individual differences (Sandri et al., 2023) and underrepresented perspectives (van der Meer et al., 2024).
Modern NLU approaches are commonly trained and tested on annotated datasets. In a subjective task, each data sample is typically labeled by a set of annotators, and differences in annotation are reconciled through aggregation techniques (e.g., majority voting), resulting in a single “gold label” (Uma et al., 2021). This approach, though effective for training ML algorithms, neglects the labels of minorities, which becomes problematic, especially, in the case of sensitive subjective tasks.
@@
Subjectivity has been addressed by modeling the full distribution of annotations for each data sample as opposed to aggregating them (Plank, 2022). However, resources for such approaches are scarce, as most datasets do not (yet) make fine-grained annotation details available (Cabitza et al., 2023), and representing a full range of perspectives is contingent on obtaining annotations from a diverse crowd (Bakker et al., 2022).
@@
One way of accounting for a limited annotation budget is to use Active Learning (Settles, 2012, AL). Given a pool of unannotated data samples, AL employs a sample selection strategy to select maximally informative samples for training, retrieving the corresponding annotations from a ground truth oracle (e.g., a single human expert). However, in subjective tasks there is no such oracle, instead we rely on a set of available annotators. Given this practical constraint, we argue that informativeness for AL manifests in both samples and annotations, as the model should also be guided to reflect the distribution of annotations. Demanding all available annotators to annotate all selected samples would provide a truthful representation of the annotation distribution, but is often unfeasible, especially if the pool of annotators is large. Thus, deciding which annotator(s) should annotate the selected samples is as critical as selecting which samples to annotate.
We introduce Annotator-Centric Active Learning (ACAL) to account for annotation diversity in subjective tasks. In ACAL, the sample selection strategy of traditional AL is followed by an annotator selection strategy as Figure 1 shows. For each data sample selected through the sample selection strategy, the annotator selection strategy selects an annotator from the available annotators. We make the following contributions: (1) We create ACAL, extending the AL approach to optimize for diversity among annotators when learning soft labels in subjective tasks. (2) We introduce a suite of annotator-centric evaluation metrics to measure both representativeness and diversity. (3) We demonstrate our approach’s effectiveness on three diverse datasets with subjective tasks—hate speech detection, moral value classification, and safety judgments.
Our experiments show that ACAL works to better approximate the distribution of human judgments with a lower annotation budget. However, this effectiveness requires a large pool of diverse annotators, as is the case for one of our datasets. In other cases, the differences between ACAL and traditional AL become smaller. Through our annotator-centric evaluation, we show that task agreement and the number of available annotations both influence the effectiveness of ACAL, hinting at a direct trade-off between learning from a majority versus being sensitive to minority annotations.
@@ 2 Related work
We review related works on annotator disagreement and active learning. Our work is novel in combining these fields to (1) represent annotation distributions through soft labels, (2) incorporate annotator selection strategies in the active learning loop, and (3) evaluate with annotator-centric metrics next to traditional evaluation.
@@ 2.1 Learning with annotator disagreement
Modeling annotator disagreement is garnering increasing attention (Aroyo and Welty, 2015; Uma et al., 2021; Plank, 2022; Cabitza et al., 2023). For instance, some aggregation methods can lead to a fairer representation than simple majority (Hovy et al., 2013; Tao et al., 2018). Alternatively, the full annotation distribution can be modeled using soft labels (e.g., Peterson et al., 2019; Müller et al., 2019; Fornaciari et al., 2021; Collins et al., 2022). Other approaches leverage annotator-specific information, e.g., by including individual classification heads per annotator (Davani et al., 2022), embedding annotator-specific behavior (Mokhberian et al., 2023), or encoding the annotator’s socio-demographic information (Beck et al., 2023).
Yet, representing annotator diversity remains challenging. Standard calibration metrics under human label variation may be unsuitable, especially when the variation is high (Baan et al., 2022). Trade-offs ought to be made between collecting more samples or more annotations Gruber et al. (2024). Further, solely measuring differences among sociodemographic traits is not sufficient to fully capture opinion diversity (Orlikowski et al., 2023). To this end, we represent diversity based on which annotators have annotated, what they annotated, and how they have annotated. We experiment with different annotator selection strategies to reveal what aspects impact task performance and annotation budget.
@@ 2.2 Active Learning
AL enables a supervised learning model to achieve high performance with a few training examples if chosen judiciously Settles (2012). In a typical AL scenario, a vast collection of unlabeled data is available, and an oracle (e.g., a human expert) can be asked to annotate this unlabeled data. A sampling strategy is employed to iteratively (and smartly) select the next batch of unlabeled data for annotation by the oracle Ren et al. (2021).
AL has found widespread application in the field of NLP Zhang et al. (2022). Two main strategies are employed, either by selecting the unlabeled samples on which the model prediction is most uncertain (Zhang et al., 2017), or by selecting samples that are most representative of the unlabeled dataset (Erdmann et al., 2019; Zhao et al., 2020).
The combination of AL and annotator diversity is a novel direction that has not garnered much attention yet. Existing work proposes to align model uncertainty with annotator uncertainty (Baumler et al., 2023), whereas others adapt annotator-specific classification heads in AL settings (Wang and Plank, 2023), or select texts to annotate based on annotator preferences (Kanclerz et al., 2023).
Existing methods ignore a crucial part of learning with human variation: the diversity among annotators. We focus on which annotators should annotate, such that it best informs us about the underlying label diversity.
@@ 3.1 Soft-Label prediction
Consider a dataset composed of triples , where is a data sample (i.e., a piece of text) and is the class label assigned by annotator . The multiple labels assigned to a sample by the different annotators are usually combined into an aggregated label . For training with soft labels, the aggregation typically takes the form of maximum likelihood estimation Uma et al. (2021):
@@ 3.2 Active Learning
AL imposes a sampling technique for inputs , such that the most informative sample(s) are picked for learning. In a typical AL approach, a set of unlabelled data points is available. At every iteration, a sample selection strategy selects samples to be annotated by an oracle that provides the ground truth label distribution . The selected samples and annotations are added to the labeled data , with which the model is trained. Alg. 1 provides an overview of the procedure. In our sample selection strategies, a batch of data of a given size is queried at each iteration. In our experiments, we compare the following strategies:
@@
predicts a distribution over class labels with for each . Select the samples with the highest prediction entropy (i.e., the samples on which the model is most uncertain).
@@ 3.3 Annotator-Centric Active Learning
The ACAL approach builds on the AL approach. In contrast to AL, which retrieves an aggregated annotation , ACAL employs an annotator selection strategy to select one annotator and their annotation for each selected data point . Alg. 2 describes the ACAL approach.
Dataset | Task (dimension) | Num. Samples | Num. Annotators | Num. Annotations | Avg. Annotations per item |
DICES | Safety Judgment | 990 | 172 | 72,103 | 72.83 |
MFTC | Morality (care) | 8434 | 23 | 31310 | 3.71 |
MFTC | Morality (loyalty) | 3288 | 23 | 12803 | 3.89 |
MFTC | Morality (betrayal) | 12546 | 23 | 47002 | 3.75 |
MHS | Hate Speech (dehumanize, genocide, respect) | 17282 | 7807 | 57980 | 3.35 |
We propose annotator selection strategies that include annotations from diverse annotators. The strategies vary in the type of information used to represent differences between annotators, and include what or how the annotators have annotated thus far. We test the following strategies:
@@
selects one random annotator .
@@
considers only the labels that annotators have assigned. Given a new sample , selects the available annotator that has the largest bias toward the minority label compared to the other available annotators, i.e., who has annotated other samples with the minority label the most. The minority label is selected as the class with the smallest annotation count in the available dataset thus far.
@@ 4.1 Datasets
Table 1 introduces the three datasets that we use, with variation in domain, annotation task (in italics), annotator count, and annotations per instance.
The DICES Corpus (Aroyo et al., 2023) is composed of 990 conversations with an LLM where 172 annotators provided judgments on whether a generated response can be deemed safe (3-way judgments: yes, no, unsure). We perform a multi-class classification with the scores.
The MFTC Corpus Hoover et al. (2020) is composed of 35K tweets that 23 annotators annotated with any of the 10 moral elements from the Moral Foundation Theory Graham et al. (2013). We select the elements of loyalty (lowest annotation count), care (medium count), and betrayal (highest count) and perform three binary classifications to predict the presence of the respective elements. As most tweets were labeled non-moral (i.e., with no moral element), we balanced the datasets by subsampling the non-moral class.
The MHS Corpus (Sachdeva et al., 2022) consists of 50K social media comments on which 8K annotators judged three hate speech aspects—dehumanize (low inter-rater agreement), respect (medium agreement), and genocide (high agreement)—on a 5-point Likert scale. We perform a multi-class classification with the annotated Likert scores for each task.
The datasets and tasks differ in the entropy scores over annotations (Appendix A.5). DICES and MHS generally have medium normalized entropy scores (most lie between , whereas the MFTC entropy scores are highly polarized.
4.2 Training procedure
We test the annotator selection strategies proposed in Section 3.3 by comparing all possible combinations of the two different sample selection strategies ( and ) with the annotator selection strategies (, , , and ). At each round, we use to select unique samples from the unlabeled data pool . We empirically select to be the smallest between 5% of the number of available annotations and the number of unique samples in the training set. For each selected sample , we use to select one annotator and retrieve their annotation .
To populate the annotation history for the annotation selection strategies, we perform a single warmup round with randomly selected annotations before starting the ACAL iterations (Zhang et al., 2022). We report our training progress results on a validation set with 3-fold cross-validation, showing the average to account for stability across random data splits (into 80% train, 10% validation, and 10% test) and initialization. Then, we select the model iteration that led to the best performance (according to ) on the validation set and evaluate it using a separate test set.
We compare our work with traditional Oracle-based AL approaches ( and ), which use the data sampling strategies but obtain all possible annotations for each sample (following Alg. 1). Moreover, we employ a passive learning (PL) approach as an upper bound by training the model on the full dataset, thus observing all available samples and annotations. Our baselines follow the analogous cross-validation setup.
@@ 4.3 Evaluation metrics
The ACAL strategies aim to guide the algorithm to model a representative distribution of the annotator’s perspectives while reducing human annotation effort. To this end, we evaluate the model both with a traditional evaluation metric and a metric aimed at comparing predicted and annotated distributions:
@@
The measures the divergence between the distribution of label annotation and prediction (Nie et al., 2020). We report the average for the samples in the test set to measure how well the algorithm can model the annotation distribution.
Next, since our proposed annotator selection strategies aim to promote diversity, we introduce novel annotator-centric evaluation metrics. First, we report the average among annotators. Second, in line with Rawls’ principle of maximum fairness Rawls (1973), the result for the worst-off annotators:
- Per-annotator ()
We compute the for each annotator in the test set using their annotations as golden labels, and average it.
- Per-annotator ()
We compute the for each annotator in the test set using their annotations as target distribution, and average it.
- Worst per-annotator ()
We compute the for each annotator in the test set using their annotations as golden labels, and report the average of the lowest 10% (to mitigate noise).
- Worst per-annotator ()
We compute the for each annotator in the test set using their annotations as target distribution, and report the average of the lowest 10% (to mitigate noise).
These evaluation metrics allow us to measure the trade-offs between modeling the majority agreement, a representative distribution of annotations, and accounting for minority voices. We report these metrics on the validation set (as progress over the AL iterations) and test set (by using the best-performing model on the validation set), as described in Section 4.2.

### v2

@@ Abstract
Active Learning (AL) addresses the high costs of collecting human annotations by strategically annotating the most informative samples. However, for subjective NLP tasks, incorporating a wide range of perspectives in the annotation process is crucial to capture the variability in human judgments. We introduce Annotator-Centric Active Learning (ACAL), which incorporates an annotator selection strategy following data sampling. Our objective is two-fold: (1) to efficiently approximate the full diversity of human judgments, and (2) to assess model performance using annotator-centric metrics, which emphasize minority perspectives over a majority. We experiment with multiple annotator selection strategies across seven subjective NLP tasks, employing both traditional and novel, human-centered evaluation metrics. Our findings indicate that ACAL improves data efficiency and excels in annotator-centric performance evaluations. However, its success depends on the availability of a sufficiently large and diverse pool of annotators to sample from.
@@ 1 Introduction
A challenging aspect of natural language understanding (NLU) is the variability of human judgment and interpretation in subjective tasks (e.g., hate speech detection) (Plank, 2022). In a subjective task, a data sample is typically labeled by a set of annotators, and differences in annotation are reconciled via majority voting, resulting in a single (supposedly, true) “gold label” (Uma et al., 2021). However, this approach has been criticized for treating label variation exclusively as noise, which is especially problematic in sensitive subjective tasks (Aroyo and Welty, 2015) since it can lead to exclusion of minority voices (Leonardelli et al., 2021).
@@
Subjectivity can be addressed by modeling the full distribution of annotations for each data sample instead of employing gold labels (Plank, 2022). However, resources for such approaches are scarce, as most datasets do not (yet) make fine-grained annotation details available (Cabitza et al., 2023), and representing a full range of perspectives is contingent on obtaining costly annotations from a diverse set of annotators (Bakker et al., 2022).
@@
One way to handle a limited annotation budget is to use Active Learning (Settles, 2012, AL). Given a pool of unannotated data samples, AL employs a sample selection strategy to obtain maximally informative samples, retrieving the corresponding annotations from a ground truth oracle (e.g., a single human expert). However, in subjective tasks, there is no such oracle. Instead, we rely on a set of available annotators. Demanding all available annotators to annotate all samples would provide a truthful representation of the annotation distribution, but is often unfeasible, especially if the pool of annotators is large. Thus, deciding which annotator(s) should annotate is as critical as deciding which samples to annotate.
In most practical applications, annotators are randomly selected. This results in an annotation distribution insensitive to outlier annotators —most annotations reflect the majority voices and fewer reflect the minority voices. This may not be desirable in applications such as hate speech, where the opinion of majority and minority should be valued equally. In such cases, a more deliberate annotator selection is required. To ensure a balanced representation of majority and minority voices, we can leverage strategies inspired by Rawls’ principle of fairness Rawls (1973), which advocates that a fair society is achieved when the well-being of the worst-off members of society (the minority annotators, in this case) is maximized.
We introduce Annotator-Centric Active Learning (ACAL) to emphasize and control who annotates which sample. In ACAL (Figure 1), the sample selection strategy of traditional AL is followed by an annotator selection strategy, indicating which of the available annotators should annotate each selected data sample.
Contributions
Our experiments show that the proposed ACAL methods can approximate the distribution of human judgments similar to AL while requiring a lower annotation budget and modeling individual and minority voices more accurately. However, our evaluation shows how the task’s annotator agreement and the number of available annotations impact ACAL’s effectiveness—ACAL is most effective when a large pool of diverse annotators is available. Importantly, our experiments show how the ACAL framework controls how models learn to represent majority and minority annotations, which is crucial for subjective and sensitive applications.
@@ 2.1 Learning with annotator disagreement
Modeling annotator disagreement is garnering increasing attention (Aroyo and Welty, 2015; Uma et al., 2021; Plank, 2022; Cabitza et al., 2023). Changing annotation aggregation methods can lead to a fairer representation than simple majority (Hovy et al., 2013; Tao et al., 2018). Alternatively, the full annotation distribution can be modeled using soft labels (Peterson et al., 2019; Müller et al., 2019; Collins et al., 2022). Other approaches leverage annotator-specific information, e.g., by including individual classification heads per annotator (Davani et al., 2022), embedding annotator behavior (Mokhberian et al., 2023), or encoding the annotator’s socio-demographic information (Beck et al., 2023). Representing annotator diversity remains challenging. Standard calibration metrics under human label variation may be unsuitable, especially when the variation is high (Baan et al., 2022). Trade-offs ought to be made between collecting more samples or more annotations Gruber et al. (2024). Further, solely measuring differences among sociodemographic traits is not sufficient to capture opinion diversity (Orlikowski et al., 2023).
We represent diversity based on which annotators annotated what and how. We experiment with annotator selection strategies to reveal what aspects impact task performance and annotation budget.
@@ 2.2 Active Learning
AL enables a supervised learning model to achieve high performance by judiciously choosing a few training examples Settles (2012). In a typical AL scenario, a large collection of unlabeled data is available, and an oracle (e.g., a human expert) is asked to annotate this unlabeled data. A sampling strategy is used to iteratively select the next batch of unlabeled data for annotation Ren et al. (2021).
AL has found widespread application in NLP Zhang et al. (2022). Two main strategies are employed, either by selecting the unlabeled samples on which the model prediction is most uncertain (Zhang et al., 2017), or by selecting samples that are most representative of the unlabeled dataset (Erdmann et al., 2019; Zhao et al., 2020).
The combination of AL and annotator diversity is a novel direction. Existing works propose to align model and annotator uncertainties (Baumler et al., 2023), adapt annotator-specific classification heads in AL settings (Wang and Plank, 2023), or select texts to annotate based on annotator preferences (Kanclerz et al., 2023). These methods ignore a crucial part of learning with human variation: the diversity among annotators. We focus on selecting annotators such that they best inform us about the underlying label diversity.
@@ 3.1 Soft-label prediction
Consider a dataset of triples , where is a data sample (i.e., a piece of text) and is the class label assigned by annotator . The multiple labels assigned to a sample by the different annotators are usually combined into an aggregated label . For training with soft labels, the aggregation typically takes the form of maximum likelihood estimation Uma et al. (2021):
@@ 3.2 Active Learning
AL imposes a sampling technique for inputs , such that the most informative sample(s) are picked for learning. In a typical AL approach, a set of unlabelled data points is available. At every iteration, a sample selection strategy selects samples to be annotated by an oracle that provides the ground truth label distribution . The selected samples and annotations are added to the labeled data , with which the model is trained. Alg. 1 provides an overview of the procedure.
In the sample selection strategies, a batch of data of a given size is queried at each iteration. Our experiments compare the following strategies:
@@
predicts a distribution over class labels with for each , and selects samples with the highest prediction entropy (the samples the model is most uncertain about).
@@ 3.3 Annotator-Centric Active Learning
ACAL builds on AL. In contrast to AL, which retrieves an aggregated annotation , ACAL employs an annotator selection strategy to select one annotator and their annotation for each selected data point . Alg. 2 describes the ACAL approach.
We propose three annotator selection strategies to gather a distribution that uniformly contains all possible (majority and minority) labels, inspired by Rawls’ principle of fairness Rawls (1973). The strategies vary in the type of information used to represent differences between annotators, including what or how the annotators have annotated thus far. Our experiments compare the following strategies:
@@
randomly selects an annotator .
@@
considers only the labels that annotators have assigned. The minority label is selected as the class with the smallest annotation count in the available dataset thus far. Given a new sample , selects the available annotator that has the largest bias toward the minority label compared to the other available annotators, i.e., who has annotated other samples with the minority label the most.
@@ 4.1 Datasets
We use three datasets which vary in domain, annotation task (in italics), annotator count, and annotations per instance.
The DICES Corpus (Aroyo et al., 2023) is composed of 990 conversations with an LLM where 172 annotators provided judgments on whether a generated response can be deemed safe (3-way judgments: yes, no, unsure). Samples have 73 annotations on average. We perform a multi-class classification with the scores.
The MFTC Corpus Hoover et al. (2020) is composed of 35K tweets that 23 annotators annotated with any of the 10 moral elements from the Moral Foundation Theory Graham et al. (2013). We select the elements of loyalty (lowest annotation count), care (average count), and betrayal (highest count). Samples have 4 annotations on average. We create three binary classifications to predict the presence of the respective elements. As most tweets were labeled as non-moral (i.e., with no moral element), we balanced the datasets by subsampling the non-moral class.
The MHS Corpus (Sachdeva et al., 2022) consists of 50K social media comments on which 8K annotators judged three hate speech aspects—dehumanize (low inter-rater agreement), respect (medium agreement), and genocide (high agreement)—on a 5-point Likert scale. Samples have 3 annotations on average. We perform a multi-class classification with the annotated Likert scores for each task.
The datasets and tasks differ in levels of annotator agreement, measured via entropy of the annotation distribution. DICES and MHS generally have medium entropy scores, whereas the MFTC entropy is highly polarized (divided between samples with very high and very low agreement). Appendix A.5 provides details of the entropy scores.
@@ 4.2 Evaluation metrics
The ACAL strategies aim to guide the model to learn a representative distribution of the annotator’s perspectives while reducing annotation effort. To this end, we evaluate the model both with a traditional evaluation metric and a metric aimed at comparing predicted and annotated distributions:
@@
The measures the divergence between the distribution of label annotation and prediction (Nie et al., 2020). We report the average for the samples in the test set to measure how well the model can represent the annotation distribution.
Further, since ACAL shifts the focus to annotators, we introduce novel annotator-centric evaluation metrics. First, we report the average among annotators. Second, in line with Rawls’ principle of fairness, the result for the worst-off annotators:
- Per-annotator () and ()
We compute the (or ) for each annotator in the test set using their annotations as golden labels (or target distribution), and average it.
- Worst per-annotator () and ()
We compute the (or ) for each annotator in the test set using their annotations as golden labels (or target distribution), and report the average of the lowest 10% (to mitigate noise).
These metrics allow us to measure the trade-offs between modeling the majority agreement, a representative distribution of annotations, and accounting for minority voices. In the next section, we describe how we obtained the results.
4.3 Training procedure
We test the annotator selection strategies proposed in Section 3.3 by comparing all combinations of the two sample selection strategies ( and ) and the four annotator selection strategies (, , , and ). At each iteration, we use to select unique samples from the unlabeled data pool . We select as the smallest between 5% of the number of available annotations and the number of unique samples in the training set. For each selected sample , we use to select one annotator and retrieve their annotation .
We split each dataset into 80% train, 10% validation, and 10% test. We start the training procedure with a warmup iteration of randomly selected annotations (Zhang et al., 2022). We proceed with the ACAL iterations by combining and . We select the model iteration that led to the best performance on the validation set and evaluate it on the test set. We repeat this process across three data splits and model initializations. We report the average scores on the test set. Appendix A contains additional details on training.
We compare ACAL with traditional oracle-based AL approaches ( and ), which use the data sampling strategies but obtain all possible annotations for each sample as in Alg. 1. Further, we employ a passive learning (PL) approach as an upper bound by training the model on the full dataset, thus observing all available samples and annotations. Similar to ACAL, the AL and PL baselines are averaged over three seeds.